40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5628     Entry time: Fri Oct 7 11:45:24 2011
Author: Koji 
Type: Summary 
Category: LSC 
Subject: POY11 installed, 55MHz PD at POY removed 

POY11 PD was installed last night. The lock of the Y arm was confirmed with the POY11I signal.

- The DC transimpedance was modified to be 1010V/A as the incident power is tiny.

- The demodulation phase of the roughly adjusted (148deg) to have PDH signal at the I-phase.
 
The comparison with AS55I signal exhibits that POY11I have ~150 times weaker signal with 45dB whitening.
 
(In total 25000 times weaker.)

On the way to make POY11 functioning, there were many fixes at the LSC rack...


Details:

- The PD interface cards (power supply for the RFPDs) were checked:
So far the two card at the right hand side were checked. 

Desipite the previous entry reported the issues on those boards, they did not show any problem yesterday.

One hypothetical possibility is the enabling switches that is controlled from the old slow epics targets.

- POY55 was removed

This 55MHz PD is supposed to be installed at POP.
The PD, an RF cable, an RF amp, the power supply of the RF amp were removed.

- POY11 was installed

The PD was placed where the 55MHz was placed.

The beam was aligned on the diode using the IR viewer and the digital multimeter.

The power supply cable and the RF cable for POY on the ITMY table were used.

There were an ND filter on the POY beam path. It was removed.


- On the LSC rack
The PD RF was connected to the patch panel at the top of the rack.

There were loose connectors on the patch panel. Some connectors were tightened on the panel.

I found that POY11 and POX11 had I&Q signal reversely connected to the whitening board.
   ==> These were fixed but
require the orthogonality test again for those channels.



The I phase output of the AS11 demod board had a broken connector. 

The onboard SMA has got disintegrated because of too much twist on the connector.
The board was once removed from the rack and the connector was fixed using a heat gun and soldering.

The DC signals were checked. POYDC was not correctly connected. POYDC were correctly connected to the POYDC channel.

- CDS
c1lsc was found with the RFM frozen.
The c1lsc machine was soft-rebooted after stopping all of the RT processes.

Once the RT processes came back, they were all burtrestored.

- PDH locking
Restored Y-arm. Locked it with AS55Q.
Ran ASS alignment for Y-arm.
100cnt 150Hz sinusoidal signal is applied to ETMY

Measured the PSD of AS55Q, POY11I, and POY11Q.
Adjusted the demod phase so that the excitation could be minimized in POY11Q.


ELOG V3.1.3-